OpenStudy (anonymous):

An ideal (non-viscous) liquid with density u is poured into a cyndrical vessel with a cross-sectional area of A1 to a level at a height h from the bottom. The bottom has an opening with a cross-sectional area A2. Find the time it takes the liquid to flow out.

OpenStudy (anonymous):

Do you have any specific question ?

OpenStudy (anonymous):

Use Bernoulli's equation (treating it as a streamline flow) : \[P + \rho*g*h + \frac{ 1 }{ 2}*\rho*v ^{2} = constant\] Apply the equation at the height h of the liquid and at the orifice of area A2; You'll get 'v' from it. A2*v= volume flow rate, i.e. volume of fluid leaving the container As the total volume of liquid in the container = A1*h Divide the two to get the time it takes for the total liquid to flow out.

OpenStudy (anonymous):

I still am not understanding this. Could you go into greater detail of their solution